OpenStudy (anonymous):

HELP!! Use the point-slope form linear equation: y+4=-4(x-2) What is the equation in standard form of a parallel line that passes through (0,-2)?

OpenStudy (mathmate):

Equation of a line in point-slope form: \((y-y_1)=m(x-x_1)\) where m=slope, \((x_1,y_1)\) is a point through which the line passes.

OpenStudy (anonymous):

sorry Im still lost

OpenStudy (mathmate):

Can you find the slope of the given line (which is also the slope of the line you are looking for)?

OpenStudy (anonymous):

I need help. Im lost :/

OpenStudy (anonymous):

I believe the slope is -4

OpenStudy (mathmate):

That is correct, and this corresponds to "m" in the point-slope form of the formula. Now you will have to substitute \(x_1,y_1\) = (0,-2) into the point slope form of the equation.

OpenStudy (anonymous):

can u set it up for me?

OpenStudy (mathmate):

It's already done in my the equation of my first post. You will have to substitute m=-4 \(x_1\)=0, \(y_1\)=-2 into the equation and simplify as you wish.

OpenStudy (mathmate):

You can post your result if you're not sure. I'm sure someone will be glad to look it over and check it for you, as I am logging off to grab a bite! :(

OpenStudy (anonymous):

What is the equation in standard form of a parallel line

OpenStudy (anonymous):

@sourwing can u help me plez

OpenStudy (anonymous):

Heres wat I have so far if its correct.. y-1-2=4(x-0) y+2=4x

OpenStudy (mathmate):

Not exactly, it should read: \(y-(-2)=-4(x-0)\) which translates to: \(y+2=-4x)\) or \(y=-4x-2)\)